OpenStudy (anonymous):

I wrote my own scenario but cant figure out how much each watermelon and yam cost. Someone please help me :) Heres my scenario: I went to the store and bought 2Xi gua (chinese watermelon) and 1Yam and spent a total $9.00. Then the next day I had bought 3Xi gua (chinese watermelon) and 4Yams and spent $26.00. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (goformit100):

2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

You have to use the test value to figure out this answer. The test value will be 0 For this example use x=0 solve then use y = 0 2(0) + y = 9 Y=9 2x + (0) = 9 /2 /2 x = 4.5 Now graph these points (4.5,9) You can do the last part.

OpenStudy (anonymous):

okay thank you soo much!!

OpenStudy (anonymous):

You are welcome.